OpenStudy (nastech):

Solve the set of differential equations using: 1) Euler's method. 2)4th order Runge Kutta method assuming at t=0, y1=4, y2=6 integrate to t=2 with step size of 0.5 dy1/dt=-0.5y1 dy2/dt=4-0.3y1-0.1y2

OpenStudy (helpppp):

@mathmate

OpenStudy (kainui):

not a question

OpenStudy (mathmate):

@nastech When are you posting the set of DE?

OpenStudy (nastech):

sorry I've edited the question

OpenStudy (loser66):

\[X=(x,y)=\left[\begin{matrix}0-0.5x+0y\\4-0.3x-0.1y\end{matrix}\right]\] \[X(0)=\left[\begin{matrix}4\\6\end{matrix}\right]\] Euler method step1 with \(\triangle t=0.5\) \[X(0.5) =\left[\begin{matrix}4\\6\end{matrix}\right]+ 0.5*\left[\begin{matrix}0-0.5*4-0*6\\4-0.3*4-0.1*6\end{matrix}\right]=\left[\begin{matrix}3\\7.1\end{matrix}\right]\]

OpenStudy (loser66):

Euler method step2 starts from step1 is \[X(0.5) =\left[\begin{matrix}3\\7.1\end{matrix}\right]+ 0.5*\left[\begin{matrix}0-0.5*3-0*7.1\\4-0.3*3-0.1*7.1\end{matrix}\right]=\left[\begin{matrix}2.25\\8.295\end{matrix}\right]\]

OpenStudy (loser66):

You do \(X(1.5),X(2)\)

OpenStudy (loser66):

Oh, I do it by hand, not sure what you want me to do. Matlab??